OpenStudy (anonymous):

Solving Quadratic Equations by Factoring. Solve each equation by factoring. 1) (k+1)(k-5)=0

jimthompson5910 (jim_thompson5910):

This has already been factored. All we need to do now is use the zero product property (k+1)(k-5)=0 k+1=0 or k-5=0 k=-1 or k=5

OpenStudy (anonymous):

What about one that doesn't have a zero? xsquared-11x+19=-5

jimthompson5910 (jim_thompson5910):

x^2-11x+19=-5 x^2-11x+19+5=0 x^2-11x+24=0 (x-8)(x-3)=0 x-8=0 or x-3=0 x=8 or x=3

OpenStudy (anonymous):

I'm going to try the next one on my own using your answer as a reference. Please correct me if I did something wrong. n^2+7n+15=5 n^2+7n+15+5=0 Is it like this so far? Or would the 5 be a negative 5?

jimthompson5910 (jim_thompson5910):

on the right side, you're basically saying 0+5, so you need to subtract it from both sides

OpenStudy (anonymous):

So, n^2+7n+10=0?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Yay! Okay so, n^2+7n+10=0 (n-5) (n-2)=0 n=5 or n=2 ?

jimthompson5910 (jim_thompson5910):

close...

jimthompson5910 (jim_thompson5910):

but it should be (n+5)(n+2)=0

jimthompson5910 (jim_thompson5910):

since 5+2 = 7 and 5*2 = 10

OpenStudy (anonymous):

But the answer has to be =-5 or =-2

jimthompson5910 (jim_thompson5910):

n^2+7n+10=0 (n+5)(n+2)=0 n+5=0 or n+2=0 n=-5 or n=-2

OpenStudy (anonymous):

Ok, got it. That makes more sense. Next one.... n^2-10n+22=-2 n^2-10n+24=0 So far, so good?

jimthompson5910 (jim_thompson5910):

yep you added 2 to both sides, so far so good

jimthompson5910 (jim_thompson5910):

now you need to find two numbers that satisfy the following a) they multiply to 24 (the last number) AND b) they add to -10 (the second coefficient)

OpenStudy (anonymous):

n^2-10n+24=0 (n+2)(n-12) Like that? Or it can't be a negative cause it's positive 24?

jimthompson5910 (jim_thompson5910):

well 2+(-12) = -10 (so far so good) but... 2*(-12) = -24, which is the opposite of what we want (positive 24)

jimthompson5910 (jim_thompson5910):

it turns out that these two numbers are: -6 and -4 Since -6+(-4) = -10 and (-6)*(-4) = 24

OpenStudy (anonymous):

Wow!! I don't even know why I hadn't thought of those two numbers. So n=-6 or n=-4

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

those two numbers help us factor, they aren't the answers

jimthompson5910 (jim_thompson5910):

n^2-10n+24=0 (n-6)(n-4)=0 ... notice the factorization has -6 and -4 (our two numbers) n-6=0 or n-4=0 n=6 or n=4 So the answers are n=6 or n=4

jimthompson5910 (jim_thompson5910):

so the answers are the opposite of the numbers that helped us factor (since we're isolating them on the other side)

OpenStudy (anonymous):

Alright, I think I'm getting it now but I'm going to try the next one since it has a positive answer. n^2+3n-12=6 n^2+3n-18=0 So I use -3 and 6, right?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

n^2+3n-12=6 n^2+3n-18=0 (n-3)(n+6)=0 n-3=0 or n+6=0 Not sure if I did that right

jimthompson5910 (jim_thompson5910):

you did that right you need to do one more step and solve for n in each equation

OpenStudy (anonymous):

Would it be turned to opposites? n=3 or n=-6

jimthompson5910 (jim_thompson5910):

yep, you got it

OpenStudy (anonymous):

Ok, so the next one is a little different. 6n^2-18n-18=6

jimthompson5910 (jim_thompson5910):

6n^2-18n-18=6 6n^2-18n-18-6=0 6n^2-18n-24=0 6(n^2-3n-4)=0 n^2-3n-4=0/6 n^2-3n-4=0 (n-4)(n+1)=0 n-4=0 or n+1=0 n=4 or n=-1

OpenStudy (anonymous):

n^2-3n-4=0/6 that step is being divided?

jimthompson5910 (jim_thompson5910):

yes, I'm dividing both sides by 6 this is so I can effectively eliminate it and not worry about it

OpenStudy (anonymous):

Ok, got it. But I can't add anything to both sides on the next question. What do I do? 7r^2-14r=-7

jimthompson5910 (jim_thompson5910):

you can add 7 to both sides

jimthompson5910 (jim_thompson5910):

why add 7? because on the right it becomes -7+7 which becomes 0

OpenStudy (anonymous):

Add 7 to -14r?

jimthompson5910 (jim_thompson5910):

yes, but you cannot combine the terms so you just leave them as they are

jimthompson5910 (jim_thompson5910):

so the left side will be 7r^2-14r+7

OpenStudy (anonymous):

Oh okay. 7r^2-14r+7=0 What happens on the next step?

jimthompson5910 (jim_thompson5910):

now factor out the GCF 7 to get 7(r^2-2r+1)=0

OpenStudy (anonymous):

Next step would be r^2-2r+1=0/7 ?

jimthompson5910 (jim_thompson5910):

yep then you would get r^2-2r+1=0

OpenStudy (anonymous):

So r-2=0 or r+1=0 ?

jimthompson5910 (jim_thompson5910):

r^2-2r+1=0 (r-1)(r-1)=0 r-1=0 or r-1=0 r=1 or r=1 So the only answer is r=1

OpenStudy (anonymous):

Okay, got it. Next one: a^2+2a-3=0 What happens if it already equals 0?

jimthompson5910 (jim_thompson5910):

you then factor and use the zero product property to solve

OpenStudy (anonymous):

This is the next page of the worksheet and the directions say to solve each equation by completing the square. Forgot to mention that

jimthompson5910 (jim_thompson5910):

a^2+2a-3=0 a^2+2a=3 a^2+2a+1=3+1 ... Note: I'm adding half of 2 squared to both sides (a+1)^2 = 4 a+1 = sqrt(4) or a+1 = -sqrt(4) a+1= 2 or a+1 = -2 a = 1 or a = -3

OpenStudy (anonymous):

Wow this is definitely confusing. So the next one... p^2+16p-22=0 p^2+16p=22 So far, so good? Not sure where to go with the next step with the half of two squared

jimthompson5910 (jim_thompson5910):

take half of the second coefficient 16 to get 8 then square it to get 64 then add 64 to both sides

jimthompson5910 (jim_thompson5910):

we do all this to get p^2+16p+64 on the left side this is a perfect square and it factors to (p+8)^2

OpenStudy (anonymous):

What happens to the 22?

jimthompson5910 (jim_thompson5910):

you're also adding 64 to the right side So add 64 to 22 to get 86

OpenStudy (anonymous):

So p^2+16p+64=86?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

What is the next step?

jimthompson5910 (jim_thompson5910):

now factor the left side

jimthompson5910 (jim_thompson5910):

So factoring the left side gives us p^2+16p+64 (p+8)(p+8) (p+8)^2 So we now have (p+8)^2 = 86

OpenStudy (anonymous):

So then p+8=sqrt86?

jimthompson5910 (jim_thompson5910):

or p+8 = -sqrt(86), remember the plus/minus

OpenStudy (anonymous):

So then p+8=sqrt86?

OpenStudy (anonymous):

Yup, already written down

jimthompson5910 (jim_thompson5910):

oh ok

OpenStudy (anonymous):

So then p+8=sqrt86?

OpenStudy (anonymous):

so p+8 = 9.3 or p+8=-9.3 ?

jimthompson5910 (jim_thompson5910):

approximately yes

OpenStudy (anonymous):

Ok. Next page the directions says to solve each equation with the quadratic formula. m^2-5m-14=0

jimthompson5910 (jim_thompson5910):

The Quadratic formula is m = (-B +- sqrt( B^2-4AC ))/(2A) In the case of m^2-5m-14, A=1, B=-5, and C=-14 m = (-B +- sqrt( B^2-4AC ))/(2A) m = (-(-5) +- sqrt( (-5)^2-4(1)(-14) ))/(2*1) m = (5 +- sqrt( (-5)^2-4(1)(-14) ))/(2*1) m = (5 +- sqrt( 25+56 ))/(2) m = (5 +- sqrt( 81 ))/(2) m = (5 +- 9)/(2) m = (5 + 9)/(2) or m = (5 - 9)/(2) m = (14)/(2) or m = (-4)/(2) m = 7 or m = -2 So the solutions are m = 7 or m = -2

OpenStudy (anonymous):

I've been working on Algebra for 3 hours now, I think its time to take a break and move on to AP History. Know anything about Andrew Jackson? Thanks a lot for helping me though. You're really patient and a good tutor so sorry for taking up a lot of your time.

jimthompson5910 (jim_thompson5910):

yeah that's probably a good idea lol I know a bit about him, but probably not anything in great depth. What do you want to know?